"What do you think is the ideal number of children for a family to have?" A Gallup Poll asked this question of 1016 randomly chosen adults. Almost half (49%) thought two children was ideal.† We are supposing that the proportion of all adults who think that two children is ideal is p = 0.49. B) What is the probability that a sample proportion p̂ falls between 0.46 and 0.52 if the sample is an SRS of size n = 5000? (Round your answer to four decimal places.)
Answer #1

Given that! P= o-49 and n=5000 (O-46G Pく0.52) 냐 P- p 0.52-0-49 5000 5o06 PC-4:24くてく4.24) = 2 , p(Z < 4.24)-1 = 2大1.0000,-! Th